The value 'V'(in dollars per acre) of a wheat crop 'd' days after planting is given by V=12d-0.5d^2 for 20
Answer by stanbon(75887)

The value 'V'(in dollars per acre) of a wheat crop 'd' days after planting is
given by V=12d-0.5d^2 for 20 < d < 80.
Find the maximum value of an acre of wheat.
How many days after it was planted will it take to achieve this maximum value?
----------------
Maximum occurs at d = -b/2a = -12/(2*-0.5) = 12
V(d) = 12*12 - 0.5(12^2) = $72 per acre
=========================
Cheers,
Stan H.